OpenStudy (anonymous):

Use the definitions of the trigonometric ratios for a right triangle to derive a cofunction identity for tan(90° - A)

OpenStudy (anonymous):

So what's the definition of tangent in terms of sine and cosine?

OpenStudy (anonymous):

??

OpenStudy (anonymous):

\[\tan A = \frac{\sin A}{\cos A}\] And if we know that \[\sin (90º-A)=\cos A\] and that \[\cos (90º-A) = \sin A\] What does that leave us with?